Number of Food Number of Service Probabilit 0.12 0.29 0.42 0.17 laints 0.18 0.38 0.34 0.10 4.83 Refer to the informati on in th e previous exercise Find the mean and standard deviation of the to- tal number of complaints received in a week. Hav- ing reached this point, you are concerned that the numbers of food and service complaints may not be independent of each other. However, you have no information about the nature of their dependence What can you now say about the mean and standard deviation of the total number of complaints received in a week?Hi, in the question 83E(statistics for buisness and economics 8th ed) they say that you have a concern that its NOT independet? but in they answer they have just made an assumption that it is inindependent. How du you calculate the question if it is dependet but you do not know to what degree?
Answer #1

x=c(0.12,0.29,0.12,0.17)
> mean(x)
[1] 0.175
> y=c(0.180.0.38,0.34,0.10)
Error: unexpected numeric constant in "y=c(0.180.0"
> y=c(0.18,0.38,0.34,0.10)
> mean(y)
[1] 0.25
> sd(x)
[1] 0.08020806
> sd(y)
[1] 0.1321615
>